OpenStudy (anonymous):

f(x) = x2 – 3x + 2 for f(-1)?

OpenStudy (anonymous):

Just plug in -1 for all the x's.

OpenStudy (anonymous):

We don't understand how to figure the math.

OpenStudy (anonymous):

I assume you mean X^2? If so: f(-1) = (-1)^2 -3(-1) + 2 = 1 + 3 + 2 = 6 f(-1) = 6